OpenStudy (anonymous):

lim x->inf (-1)^x = ? lim x->inf ((-1)^x)*x = ? lim x->inf x = ?

OpenStudy (turingtest):

start with the easiest\[\lim_{x\to\infty}x=\]what do you think?

OpenStudy (anonymous):

infinity

OpenStudy (turingtest):

of course!

OpenStudy (experimentx):

the rest are undefined ..

OpenStudy (anonymous):

for the first on maple it comes -1-I .. 1+I

OpenStudy (turingtest):

@experimentX what you just did is called "sniping" - when I am trying to provide an understanding and you burst in with just the answer please do not do that, it lowers the value of the site

OpenStudy (turingtest):

@woundedtiger40 next most difficult (in my opinion)\[\lim_{x\to\infty}(-1)^x\]now what are some of the terms of this sequence ? what f(x)=(-1)^x what is f(0), f(1), etc...?

OpenStudy (anonymous):

@TuringTest please go on

OpenStudy (anonymous):

1, -1,1,-1,1..... it's an alternating series

OpenStudy (turingtest):

so does it ever approach one particular value?

OpenStudy (anonymous):

no OK so the limit D.N.E.

OpenStudy (turingtest):

exactly :D

OpenStudy (anonymous):

hence it's product will not exist

OpenStudy (anonymous):

am I right

OpenStudy (anonymous):

@TuringTest you rock man

OpenStudy (anonymous):

thankssssssss

OpenStudy (turingtest):

yes, because the (-1)^n part is making each point jump from one side of the x-axis to the other, AND f(x)=x does not converge to zero the limit of their product DNE -You Rock WT!

OpenStudy (anonymous):

@TuringTest I mean you are good in teaching

OpenStudy (anonymous):

explaining things...

OpenStudy (turingtest):

Thank you, I try :D -\(if\) f(x) \(did\) converge to zero though, then the limit to infty may exist for example\[\lim_{x\to\infty}{(-1)^x\over x}=0\]because even though the whole thing oscillates, the oscillation gets smaller and smaller, eventually just being a straight line (i.e. y=0) http://www.wolframalpha.com/input/?i=plot+%28-1%29%5Ex%2Fx